# Tighter Privacy Auditing of DP-SGD in the Hidden State Threat Model

### v2

@@ 5.2 Auditing results for periodicity $k=1$
Low-dimensional models. Auditing smaller models is more challenging because the ratio between the mini-batch size and the number of parameters is much larger. To evaluate our adversaries in this regime, we switch to the Housing dataset with the FCNN model, which has only 68 parameters. We report the performance of our two adversaries over five independent runs in Figure 2c. We observe that randomly selecting a dimension ($\mathcal{A}_{GC}$-R) no longer achieves tight results, although it still outperforms the baseline loss-based adversary. Remarkably, our second adversary, which simulates the training algorithm to select an appropriate dimension ($\mathcal{A}_{GC}$-S), recovers nearly tight results with small variance between runs.
@@ 5.3 Auditing Results at Periodicity $k>1$
We now study how our adversaries perform when the crafted gradient is inserted at a higher periodicity $k\in\{5,25\}$. We observe in Figure 3 that (i) we still outperform the baseline loss-based adversary $\mathcal{A}_{L}$ by a large margin, but (ii) we no longer match the privacy upper bound, and our adversary becomes weaker as we increase $k$. Intuitively, the latter is due to the accumulation of the noise added on genuine gradients during the $k-1$ iterations between each crafted gradient insertion. However, privacy accounting (i.e., the upper bound) does not take advantage of this accumulated noise: as the insertion of a crafted gradient at a given step could bias subsequent genuine gradients towards a particular direction, the best one can do is to resort to the post-processing property of DP, which ensures that subsequent genuine gradients do not weaken the privacy guarantees. In the next section, we investigate whether it is possible to match the upper bound given by privacy accounting in this setting, which amounts to designing an adversary capable of crafting gradients that maximally influence subsequent genuine gradients.
@@
Using auditing via Gaussian DP (Nasr et al., 2023) for $k>1$ implies approximating the trade-off function of the mechanism with a Gaussian one (see Appendix A for relevant definitions). This approximation can be justified by the Central Limit Theorem (Dong et al., 2022), and has been used before when auditing the hidden state (Nasr et al., 2023) We show in Appendix E (Figure 7a) that the approximation error is indeed negligible in our case.
@@ 6 Towards a Worst-Case Adversary for the Hidden State
where the first step of the process is either $\theta_{1}=Z_{1}$ (a process that did not use the crafted gradient) or $\overline{\theta}_{1}=\nabla^{*}+Z_{1}$ (a process that used the crafted gradient $\nabla^{*}$), with $Z_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2})$, and $g:\mathbb{R}\rightarrow[-BC,BC]$ is a function that abstracts away the sum of gradients of the loss on a batch of size $B$ in the DP-SGD update (Equation 2). The goal of the adversary is to choose the function $g$ and the crafted gradient $\nabla^{*}$, without knowing the intermediate updates, so as to make the distribution of $\theta_{T}$ and $\overline{\theta}_{T}$ as “distinguishable” as possible.
@@
Figure 4: Auditing performance of our adversary $\mathcal{A}^{h^{*}}_{S}$ across $T=25$ steps. Figure 4a shows the evolution of the auditing performance across time for $\sigma\in\{1,8\}$ and batch size $B\in\{1,2,4,8,16\}.$ Figure 4a gives the privacy amplification rate, i.e., the ratio $\hat{\varepsilon}_{t=25}/\hat{\varepsilon}_{t=1}$ between the privacy auditing lower bounds at step 25 $(\hat{\varepsilon}_{t=25})$ and at step 1 $(\hat{\varepsilon}_{t=1})$.
Auditing with $\mathcal{A}^{h^{*}}_{S}$. Figure 4a shows the auditing performance of our adversary $\mathcal{A}_{S}^{h^{*}}$ for $T\in\{1,\dots,25\}$ steps, batch sizes $B\in\{1,2,4,8,16\}$ and Gaussian noise variances $\sigma\in\{1,8\}$ (more values are given in Figure 14 of the appendix). [^3] As in Section 5, we need to approximate the trade-off function of our mechanism with a Gaussian one (see Remark 3). In practice, the approximation error is again negligible (see Figure 7b in Appendix E). In Figure 4b, we present the ratio between the privacy auditing lower bound $\hat{\varepsilon}_{t=25}$ at the last step and the one obtained after the first step to measure the extent of the privacy amplification phenomenon for various batch sizes $B$ and values of $\sigma$. When this ratio is equal to $1$, there is no amplification and the audit is tight, while values smaller than 1 indicate the existence of a privacy amplification effect (the smaller the value, the stronger the amplification). The general trend we observe is that (i) the amplification rate increases with $\sigma$, as higher variance helps to forget initial conditions in the pair of stochastic processes (3), and (ii) the amplification rate decreases with $B$, since a larger batch implies that subsequent updates can be larger in magnitude, enabling better propagation of the initial conditions across iterations despite the noise.
The first important takeaway is that our adversary $\mathcal{A}^{h^{*}}_{S}$ achieves a tight audit when the batch size $B$ is large enough relative to the noise variance $\sigma$. In this regime, this adversary is thus optimal and there is no privacy amplification.
@@
The second key observation is that privacy amplification seems to occur when the batch size is small enough relative to the noise variance. As seen from Figure 4a, the privacy loss of $\mathcal{A}^{h^{*}}_{S}$ converges to a constant: this is because the privacy loss distributions become sufficiently far away and stop mixing over time. This constant is smaller than the upper bound provided by privacy accounting, but remains positive, indicating that this potential amplification must be qualitatively weaker than in the convex case.
@@ 7 Discussion and Future Work
Our work allows tighter auditing of DP-SGD in the hidden state model, and our results yield several important implications that advance the understanding of this threat model. For instance, Implication 3 means that the property of converging privacy loss of DP-SGD as $T\rightarrow+\infty$ recently established by Altschuler and Talwar (2022) in the convex case cannot hold for non-convex models in the general case without additional constraints. Beyond this negative result, our work suggests that a (weaker) form of privacy amplification does occur for non-convex problems in the hidden state model when the batch size is small relative to the noise variance.
Beyond differential privacy, our results have consequences for machine unlearning (Bourtoule et al., 2021), which aims to remove a specific data point from a trained model as if the data had never been used for training. A recent approach to unlearning relies on noisy training and privacy amplification by iteration (Chien et al., 2024). Implication 3 shows that this approach cannot provide complete data point deletion when considering non-convex models like neural networks.
@@ Appendix C Details about our Adversaries $\mathcal{A}_{GC}$
           $S\leftarrow S+\theta_{i}-\theta_{i-1}$
@@
        $S\leftarrow\theta_{T}-\theta_{0}$
@@ Appendix D Experimental Details on Training & Auditing
In our experimental section, we audit three neural network architectures, a ConvNet and a ResNet on the CIFAR10 dataset and a Fully-Connected Neural Network on the Housing dataset. These models were implemented using PyTorch [Paszke et al., 2019], and the DP-SGD optimizer was Opacus [Yousefpour et al., 2021]. The hyperparameters for training, privacy guarantees, and privacy auditing parameters for each audited machine learning model are meticulously outlined in Table 2. The ConvNet and ResNet have been audited when trained for 250, 1250, and 6250 optimization steps, totalling approximately 3500 GPU hours.
For each experiment, we audit $R=5000$ machine learning models using GDP auditing (see Appendix A). For the Clopper Pearson confidence intervals, we use a confidence of $95\%$ (consistent with prior work [Jagielski et al., 2020, Nasr et al., 2021]) to get a lower bound on the Type I error rate ($\alpha)$ and Type II error rate ($\beta$) for the adversary’s performance. RejectionRule is computed by selecting the best linear threshold classifier on $S$. While, in practice, the adversary should pick the classifier on a set of held-out scores $S$, we select the classifier directly on $S$ to save the computational resources required to compute the held-out set $S$. We stress that it is not an issue, as differential privacy guarantees hold against any choice of classifier.
Table 2: Hyperparameters used per model
@@ Appendix G Additional Experimental Results
Alternative gradient-crafting strategy. The adversaries that we propose in Section 4 and use in Section 5 construct a gradient biased in a single dimension chosen by the adversary. We investigate here an alternative approach inspired by the auditing method of Andrew et al. [2024]. In their approach, the crafted gradient is randomly sampled from the unit sphere and then scaled by the sensitivity $C$. For the confidence score (RankSample in Algorithm 1), the adversary computes the cosine similarity between the canary gradient and the change in model parameters $\theta_{T}-\theta_{0}$, which is compatible with the hidden state model. For completeness, Algorithms 6-7 formally show how the generation of the gradient and the computation of the confidence score is performed. We denote this adversary by $\mathcal{A}_{CotS}$.

### v3

@@ 5.2 Auditing results for periodicity $k=1$
Low-dimensional models. Auditing smaller models is more challenging because the ratio between the mini-batch size and the number of parameters is larger. To evaluate our adversaries in this regime, we switch to the Housing dataset with the FCNN model, which has only 68 parameters. We report the performance of our two adversaries over five runs in Figure 1c. We observe that randomly selecting a dimension ($\mathcal{A}_{GC}$-R) no longer achieves tight results, although it still outperforms the baseline loss-based adversary. Remarkably, our second adversary, which simulates the training algorithm to select an appropriate dimension ($\mathcal{A}_{GC}$-S), recovers nearly tight results with small variance between runs.
@@ 5.3 Auditing Results at Periodicity $k>1$
We now study how our adversaries perform when the crafted gradient is inserted at a higher periodicity $k\in\{5,25\}$. We observe in Figure 2 that (i) we still outperform the baseline loss-based adversary $\mathcal{A}_{L}$ by a large margin, but (ii) we no longer match the privacy upper bound, and our adversary becomes weaker as we increase $k$. Intuitively, the latter is due to the accumulation of the noise added on genuine gradients during the $k-1$ iterations between each crafted gradient insertion. However, privacy accounting (i.e., the upper bound) does not take advantage of this accumulated noise: as the insertion of a crafted gradient at a given step could bias subsequent genuine gradients towards a particular direction, the best one can do is to resort to the post-processing property of DP, which ensures that subsequent genuine gradients do not weaken the privacy guarantees. In the next section, we investigate whether it is possible to match the privacy accounting upper bound in this setting, which amounts to designing an adversary capable of crafting gradients that maximally influence subsequent genuine gradients.
@@
Using auditing via Gaussian DP (Nasr et al., 2023) for $k>1$ implies approximating the trade-off function of the mechanism with a Gaussian one (see Appendix A for definitions). While this approximation can underestimate the privacy guarantees in certain cases like subsampled or shuffled mechanisms (see Dong et al. (2022); Wang et al. (2024)), in our context it is justified by the Central Limit Theorem (Dong et al., 2022), and has been used before when auditing the hidden state (Nasr et al., 2023). Appendix E (Figure 6a) shows the approximation error is indeed negligible in our case.
@@ 6 Towards a Worst-Case Adversary for the Hidden State
where the first step is either $\theta_{1}=Z_{1}$ (a process that did not use the crafted gradient) or $\overline{\theta}_{1}=\nabla^{*}+Z_{1}$ (a process that used the crafted gradient $\nabla^{*}$), with $Z_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2})$, and $g:\mathbb{R}\rightarrow[-BC,BC]$ is a function that abstracts the sum of gradients of the loss on a batch of size $B$ in the DP-SGD update (Equation 2). The goal of the adversary is to choose the function $g$ and the gradient $\nabla^{*}$, without knowing the intermediate updates, making the distribution of $\theta_{T}$ and $\overline{\theta}_{T}$ as “distinguishable” as possible.
@@
Figure 3: Auditing performance of our adversary $\mathcal{A}^{h^{*}}_{S}$ across $T=25$ steps. Figure 3a shows the evolution of the auditing performance across time for $\sigma\in\{1,8\}$ and batch size $B\in\{1,2,4,8,16\}.$ Figure 3b gives the privacy amplification rate, i.e., the ratio $\hat{\varepsilon}_{t=25}/\hat{\varepsilon}_{t=1}$ between the privacy auditing lower bounds at step 25 $(\hat{\varepsilon}_{t=25})$ and at step 1 $(\hat{\varepsilon}_{t=1})$.
Auditing with $\mathcal{A}^{h^{*}}_{S}$. Figure 3a shows the auditing performance of our adversary $\mathcal{A}_{S}^{h^{*}}$ for $T\in\{1,\dots,25\}$ steps, batch sizes $B\in\{1,2,4,8,16\}$ and noise variances $\sigma\in\{1,8\}$ (for more values see Figure 14 in the appendix). [^3] As in Section 5, we need to approximate the trade-off function of our mechanism with a Gaussian one (see Remark 3). In practice, the approximation error is again negligible (see Figure 6b in Appendix E). Figure 3b gives the ratio between the privacy auditing lower bound $\hat{\varepsilon}_{t=25}$ at the last step and the one obtained after the first step, to measure the extent of the privacy amplification phenomenon for various batch sizes $B$ and values of $\sigma$. When this ratio is equal to $1$, there is no amplification and the audit is tight, while values smaller than 1 indicate the existence of a privacy amplification effect. We observe two general trends: the amplification rate increases with $\sigma$ and decreases with $B$. This can be explained as follows. Higher noise variance ($\sigma$) helps forget initial conditions in the stochastic processes (3). Conversely, a larger batch size ($B$) allows subsequent updates to be larger, enabling the initial conditions to propagate better across iterations despite the noise. Indeed, as gradient updates are bounded by $[-BC,BC]$ and the noise remains independent of $B$, increasing $B$ allows the adversary to retain more “signal” about the canary’s presence in subsequent gradients.
Unlike the case where the crafted gradient is inserted at every step (see Implication 1), the above results reveal two distinct two regimes, as highlighted in Implications 2 and 3 below. On the one hand, when the batch size $B$ is large enough relative to the noise variance $\sigma$, our adversary $\mathcal{A}^{h^{*}}_{S}$ achieves a tight audit. In this regime, this adversary is thus optimal and there is no privacy amplification.
@@
On the other hand, privacy amplification seems to occur when the batch size is small enough relative to the noise variance. As seen from Figure 3a, the privacy loss of $\mathcal{A}^{h^{*}}_{S}$ converges to a constant: this is because the privacy loss distributions become sufficiently far away and stop mixing over time. This constant is smaller than the upper bound from privacy accounting, but remains positive, indicating that this potential amplification must be qualitatively weaker than in the convex case.
@@ 7 Discussion and Future Work
Our work allows tighter auditing of DP-SGD in the hidden state model, and our results yield several implications that advance the understanding of this threat model. For instance, Implication 3 means that the property of converging privacy loss of DP-SGD as $T\rightarrow+\infty$ recently established by Altschuler and Talwar (2022) in the convex case cannot hold for non-convex models without additional constraints. Beyond this negative result, our work suggests that a (weaker) form of privacy amplification does occur for non-convex problems when the batch size is small relative to the noise variance.
Beyond differential privacy, our results have consequences for machine unlearning (Bourtoule et al., 2021), which aims to remove a specific data point from a trained model as if it had never been used for training. A popular definition for unlearning is inspired by DP (Guo et al., 2020; Neel et al., 2021; Gupta et al., 2021), with recent work relying on noisy training and privacy amplification by iteration to provide unlearning guarantees (Chien et al., 2024). Implication 3 shows that this approach cannot completely unlearn a data point when considering non-convex models like neural networks.
@@ Appendix C Details about our Adversaries $\mathcal{A}_{GC}$
           $S\leftarrow S+\left(\theta_{i}-\theta_{i-1}\right)^{2}$
@@
        $S\leftarrow\left|\theta_{T}-\theta_{0}\right|$
     else
        $S\leftarrow\sqrt{S}$
@@ Appendix D Experimental Details on Training & Auditing
In our experimental section, we audit three neural network architectures, a ConvNet and a ResNet on the CIFAR10 dataset and a Fully-Connected Neural Network on the Housing dataset. These models were implemented using PyTorch [Paszke et al., 2019], and the DP-SGD optimizer was Opacus [Yousefpour et al., 2021]. The hyperparameters for training, privacy guarantees, and privacy auditing parameters for each audited machine learning model are meticulously outlined in Table 2. The ConvNet and ResNet have been audited when trained for 250, 1250, and 6250 optimization steps, totaling approximately 3500 GPU hours.
Algorithm 4 Privacy auditing with our gradient-crafting adversaries
  Input: dataset $D$, mini-batches $\{B_{1}\dots B_{T}\}$, periodicity $k$, number of auditing runs $R$,           privacy parameter $\delta$, noise variance $\sigma^{2}$, clipping norm $C$, initial parameters $\theta_{0}\in\mathcal{R}^{p}$,           standard Gaussian CDF $\Phi$, gradient-crafting adversary $\mathcal{A}_{GC}$
  $\nabla^{*},d\leftarrow\mathcal{A}_{GC}$triangleright get crafted gradient and biased dimension from adversary
  $S\leftarrow[]$, $b\leftarrow[]$
  for $j=1$ to $R$ do
     $b_{i}\leftarrow\text{Ber}(\frac{1}{2})$triangleright draw a random bit
     for $t\in\{1,\cdots,T\}$ do
        $\theta_{t+1}=\theta_{t}-\frac{\eta}{\left|B_{t}\right|}\left(\underset{x_{i}\in B_{t}}{\sum}\text{clip}\left(\nabla_{\theta_{t}}\ell\left(\theta_{t};x\right),C\right)+Z_{t}\right),Z_{t}\sim\mathcal{N}(0,\sigma^{2})$
        if $t\mod k=0\,\,\,\ \text{AND}\,\,\,b_{i}=0$ then
           $\theta_{t+1}=\theta_{t+1}-\frac{\eta}{|B_{t}|}\nabla^{*}$triangleright Canary is inserted
        end if
     end for
     $S_{i}\leftarrow\theta_{T}\left[d\right]-\theta_{0}[d]$triangleright Retrieve the model value at biased dimension $d$
  end for
  $S_{\text{sorted}}\leftarrow\texttt{sort}(S)$
  $H\leftarrow\Big{\{}\,h_{i}=\frac{S_{\text{sorted}}[i]+S_{\text{sorted}}[i+1]}{2}\,:\,i=1,\dots,|S_{\text{sorted}}|-1\,\Big{\}}$
  for $h_{i}\in H$ do
     $\hat{b}\leftarrow\{\mathbbm{1}\left[S_{i}\leq h_{i}\right]\;:\;i=1,\dots,|S|\}$
     TN, TP, FN, FP $\leftarrow\texttt{ConfusionMatrix}(\hat{b},b)$
     $\alpha_{i},\beta_{i}\leftarrow\texttt{ClopperPearson}(\text{TN},\text{TP},\text{FN},\text{FP})$
     $\mu_{i}\leftarrow\Phi^{-1}(1-\alpha_{i})-\Phi^{-1}(\beta_{i})$
[…]
  end for
  return $\max\{\hat{\varepsilon}_{i}:i=1,\dots,R\}$
For clarity, in Algorithm 4 we detail the Gaussian DP-based auditing procedure used for our gradient-crafting adversaries ($\mathcal{A}_{GC}$-R or $\mathcal{A}_{GC}$-S). For each experiment, we audit $R=5000$ machine learning models using GDP auditing (see Appendix A). For the Clopper Pearson confidence intervals, we use a confidence of $95\%$ (consistent with prior work [Jagielski et al., 2020, Nasr et al., 2021]) to get a lower bound on the Type I error rate ($\alpha)$ and Type II error rate ($\beta$) for the adversary’s performance. RejectionRule is computed by selecting the best linear threshold classifier on $S$. While, in practice, the adversary should pick the classifier on a set of held-out scores $S$, we select the classifier directly on $S$ to save the computational resources required to compute the held-out set $S$. We stress that it is not an issue, as differential privacy guarantees hold against any choice of classifier.
@@ Appendix G Additional Experimental Results
Alternative gradient-crafting strategy. The adversaries that we propose in Section 4 and use in Section 5 construct a gradient biased in a single dimension chosen by the adversary. We investigate here an alternative approach inspired by the auditing method of Andrew et al. [2024]. In their approach, the crafted gradient is randomly sampled from the unit sphere and then scaled by the sensitivity $C$. For the confidence score (RankSample in Algorithm 1), the adversary computes the cosine similarity between the canary gradient and the change in model parameters $\theta_{T}-\theta_{0}$, which is compatible with the hidden state model. For completeness, Algorithms 7-8 formally show how the generation of the gradient and the computation of the confidence score is performed. We denote this adversary by $\mathcal{A}_{CotS}$.